OpenStudy (hxkage):

If cosA - sinA = root 2 sinA, prove that cosA + sinA = root 2 cosA

imqwerty (imqwerty):

any ideas??

OpenStudy (hxkage):

Oh hello, Aayush bhai! :P Maybe, we could start off by squaring both the sides

imqwerty (imqwerty):

Hello :)) And yeah that will be a great start

OpenStudy (hxkage):

\[(cosA - sinA)^{2} = (\sqrt{2}sinA)^{2}\]

OpenStudy (hxkage):

Thank you! ;-;

imqwerty (imqwerty):

Try to square both equation and then add them Aw np :)

OpenStudy (hxkage):

I didn't get you there.. ;-;

imqwerty (imqwerty):

If the second equation is true then IF the RHS equals the LHS when you square and add the both the equation then we can say that second equation is true

imqwerty (imqwerty):

Theres one more way to do this tho js

OpenStudy (hxkage):

Easier?

imqwerty (imqwerty):

Yeah the other method is a bit less confusing

OpenStudy (hxkage):

Enlighten me, bro :P

imqwerty (imqwerty):

Other method- Step one: square both sides of the already given equation

OpenStudy (hxkage):

I see

OpenStudy (hxkage):

\[(cosA - sinA)^{2} = 2sinA\] ??

OpenStudy (hxkage):

Then we could expand LHS?

imqwerty (imqwerty):

Yeah expand the LHS And it must be sin^2A on the right side tho

OpenStudy (hxkage):

Oh yea, that's what I meant xD

imqwerty (imqwerty):

xD okay expand the express

OpenStudy (hxkage):

cos^2A + sin^2A - 2sinAcosA?

imqwerty (imqwerty):

Expression*

imqwerty (imqwerty):

Yeah right Now sin^2A+cos^2A is 1

OpenStudy (hxkage):

Right

imqwerty (imqwerty):

Now we will move the 2sin^2A to the left and - 2sinAcosA to the right

imqwerty (imqwerty):

So you'll get 1-2sin^A=2sinAcosA oky??

OpenStudy (hxkage):

Oh, u just exchange places.. right?

imqwerty (imqwerty):

Yeah

imqwerty (imqwerty):

Next we add 1 on both the sides

imqwerty (imqwerty):

1+1-2sin^A=1+2sinAcosA 2(1-sin^2A)=sin^2A+ cos^2A +2sinAcosA

OpenStudy (hxkage):

Oh

OpenStudy (hxkage):

Then we change sin^2A + cos^A in RHS to 1?

imqwerty (imqwerty):

Yeah the 1 on the RHS was changed into sin^2A+cos^2A

imqwerty (imqwerty):

Soo are all those steps clear?

OpenStudy (hxkage):

Yup!

imqwerty (imqwerty):

Alright :D next we write that 1-sin^2A as cos^2A andd the RHS as (cosA+sinA) ^2

OpenStudy (hxkage):

Oh

imqwerty (imqwerty):

Okay what should we do next

OpenStudy (hxkage):

Hmm

OpenStudy (hxkage):

Idk

OpenStudy (hxkage):

;-;

imqwerty (imqwerty):

Okay you're confused somewhere Tell me where

OpenStudy (hxkage):

Now we have, cos^2A = (cos+sin)^2, right?

imqwerty (imqwerty):

Theres a 2 on the left 2(cos^2A) <-LHS

OpenStudy (hxkage):

Oh

imqwerty (imqwerty):

Yeah now we just gotta square root LHS and RHS

OpenStudy (hxkage):

2cosA = cosA + sinA?

OpenStudy (hxkage):

Then we could take cosA to the left, so cosA = sinA?

imqwerty (imqwerty):

That will be root2cosA

imqwerty (imqwerty):

So you have sinA+cosA=root2cosA

OpenStudy (hxkage):

Oh yes, I'm soo careless

OpenStudy (hxkage):

Aha, its proved then!

imqwerty (imqwerty):

Yeah =]

OpenStudy (hxkage):

Tysm! :D

imqwerty (imqwerty):

Np :)